I am having problems with this problem
Find the length X in the triangle:
express in simplified radical form:
one side is 12
and the other side is 8
thank you alot I am not that good at these.

Answer by checkley75(3666)

IF THIS IS A RIGHT TRIANGLE & YOU ARE LOKING FOR THE HYPOTENUSE (THE LONGEST SIDE) THEN:
12^2+8^2=H^2
144+64=H^2
208=H^2
H=SQRT208
H=14.42